OpenStudy (anonymous):

find the slope, if it exists, of the line containing the pair of points. (5,8) and (9,-6)

OpenStudy (saifoo.khan):

\[\frac{8+6}{5-9}\]\[\frac{14}{-4}\]\[\frac{7}{-2}\]

OpenStudy (saifoo.khan):

Mine is correct.

OpenStudy (anonymous):

yea its his.